### v1

@@ 1 Introduction
Reasoning models, including ChatGPT-o1 [jaech2024openai], ChatGPT-o3 [openai2025o3], and DeepSeek-R1 [guo2025deepseek], have demonstrated remarkable capabilities across diverse and complex reasoning tasks, such as code generation, mathematical problem-solving, multi-hop question answering, and decision-making [li2025system, jiang2024survey, forootani2025survey]. A pivotal factor behind this success is the Chain-of-Thought (CoT) prompting strategy [wei2022chain], which guides models to explicitly generate intermediate reasoning steps before concluding with the final answer. This approach significantly enhances the accuracy and interpretability of the model’s outputs by effectively leveraging its intrinsic reasoning capabilities. Beyond achieving high performance, ensuring the safety of reasoning models is increasingly critical [zhou2025hidden, wang2025safety]. As these models become more prevalent in real-world applications, it is essential to prevent them from generating harmful, unethical, or adversarial content.
However, the safety mechanisms acquired by reasoning models through alignment algorithms [ouyang2022training, rafailov2023direct] demonstrate certain effectiveness in handling standard (vanilla) jailbreak prompts, but they still exhibit vulnerabilities when confronted with more complex adversarial jailbreak techniques. Figure 1, based on the MaliciousEducator dataset [kuo2025h], illustrates examples of vanilla harmful and benign prompts, along with their adversarial variants. Notably, adversarial harmful prompts frequently conceal malicious intent through indirect, covert strategies, such as framing harmful requests as educational inquiries. Conversely, adversarial benign prompts structurally resemble jailbreak attempts but do not harbor malicious intent. As demonstrated in the model responses shown in Figure 1, the model should be capable of refusing responses to adversarial harmful prompts, while avoiding excessive rejection of adversarial benign prompts.
Additionally, adversarial jailbreak prompts generally exhibit two distinctive features: (1) Limited availability, which includes a small number of prompts (usually only dozens) and an almost complete absence of corresponding reasoning responses suitable for fine-tuning. (2) High structural and semantic variability, such as mathematical encodings and pedagogically styled questions [kuo2025h, wildteaming2024, bethany2024jailbreaking]. These characteristics pose significant difficulties for traditional alignment methods reliant on extensive human intervention, limiting their scalability and responsiveness to rapidly evolving threats.
@@
We detail our methodology further in Section 3. Extensive experiments across two reasoning models and three datasets demonstrate our framework’s efficacy. Notably, on the MaliciousEducator dataset, SGASA achieved improvements of 34.5% and 32.2% using DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B [guo2025deepseek] as backbones, respectively. We also explore the impact of data volume and ratio on alignment fine-tuning results and perform cross-dataset evaluations, verifying our approach’s generalization across different attack strategies. Detailed discussions and findings are presented in Section 4.
@@ 2 Related Work
Ensuring the safety alignment of large language models has been a key research direction to guarantee their reliability and prevent the generation of harmful content. A substantial body of work has explored various alignment methods. For instance, RLHF [ouyang2022training] trains a reward model based on human rankings of multiple model outputs, and then fine-tunes the language model using this reward signal to produce responses preferred by humans. DPO [rafailov2023direct] bypasses the separate reward-model training step by converting human comparison data directly into a loss on the model’s own policy. Beyond these, with the improvement of model reasoning capabilities – especially the emergence of large reasoning models such as OpenAI’s o1 series [jaech2024openai], DeepSeek-R1 [liu2024deepseek], Alibaba’s Qwen model [Qwen3], Google’s Gemini Flash Thinking model [gemini_flash_thinking], and the KIMI reasoning model [team2025kimi] – a set of “slow thinking” alignment techniques has been proposed to enhance model safety. In this line of work, researchers directly teach the model textual safety specifications and train it to explicitly recall and reason over these rules before generating an answer. For example, deliberative alignment [guan2024deliberative] teaches the model to apply written policy rules via Chain-of-Thought reasoning. Applied to OpenAI’s o‑series models, deliberative alignment demonstrated that models can employ Chain‑of‑Thought reasoning to identify relevant policy clauses and produce safer, more compliant responses, achieving Pareto improvements by reducing over‑refusal while boosting adversarial robustness. Similarly, SafeChain [jiang2025safechain] fine-tunes models on CoT–style safety data. Other approaches address related issues, such as reducing undue refusals through multi-turn reasoning [wang2024don] or using introspective CoT analysis to identify safety risks [zhang2025safety].
@@
Many studies focus on how to bypass a model’s safety alignment mechanisms to induce it to generate harmful content, using approaches such as decoding strategies, prefilling strategies [vega2023bypassing, wei2023jailbroken], and optimization strategies [yang2023shadow, he2024your, li2025tf]. Another line of work hides harmful intent within specific contextual forms. This approach misleads the model with seemingly benign but actually malicious inputs, causing it to output unsafe information. For example, [bethany2024jailbreaking] transforms harmful user requests into mathematical symbols, while [kuo2025h] disguises them as educational scenarios (such as law enforcement training). The approach proposed in this paper mainly targets this type of jailbreak attack conducted through natural language.
@@
Given the rapid advances in the capabilities of Large Language Models (LLMs), traditional alignment approaches reliant on extensive human annotation have become increasingly difficult to scale. Consequently, a range of automated alignment methods have been extensively studied to mitigate these scalability limitations. A notable category among these methods is model self-alignment. Sun et al. (2023) [sun2023principle] proposed principle-driven self-alignment, aiming to explicitly guide models according to a concise set of human-defined principles, thus significantly minimizing the need for human supervision. Zhang et al. (2024) [zhang2024self] introduced a self-alignment strategy designed to reduce factual hallucinations by exploiting the intrinsic self-evaluation capabilities of language models. Their proposed approach, consisting of a self-evaluation mechanism combined with Self-Knowledge Tuning, demonstrated considerable improvements in factual accuracy without the need for external human annotations. Furthermore, Xu et al. (2025)  [xu2025refining] presented PT-ALIGN, a dual safety self-alignment method targeting both positive and toxic samples. By integrating Maximum Likelihood Estimation for positive examples and fine-grained Unlikelihood Training for toxic ones, PT-ALIGN effectively enhances the safety of model outputs while substantially reducing human intervention. Overall, the research on self-alignment methods has shown significant potential in reducing human involvement and improving output quality, establishing itself as an increasingly critical direction within the broader field of language model alignment [cao2024towards].
@@ 4.1 Experimental Settings
  • WildJailbreak [wildteaming2024] is an open-source safety dataset whose test set includes two types of prompts: adversarial harmful prompts, which convey harmful requests in more convoluted and stealthy ways, and adversarial benign prompts, which are similar in form but do not carry harmful intent. In our study, we construct a training subset by sampling five prompts of each type from the original training set. For evaluation, we randomly select 210 prompts of each type to create a test set comprising a total of 420 prompts.
  • MathPrompt [bethany2024jailbreaking] proposes a novel jailbreaking technique that reformulates harmful user requests into symbolic mathematical expressions to bypass safety mechanisms. Since the original dataset used in the study is not publicly available, we reconstruct the test set based on the described methodology. For the training set, we include five prompts of each type. The testing set consists of 150 harmful prompts and 150 benign prompts.
  • MaliciousEducator[kuo2025h] is a jailbreak dataset developed for educational purposes. It contains 50 samples featuring highly dangerous queries and suspicious intents. Based on this dataset, we add an equal number of benign prompts with similar structure and format. We randomly select 5 samples of each type to form the training set. The resulting test set consists of 45 harmful samples and 45 benign samples.
@@
Large Reasoning Models. In this work, we investigate two reasoning models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B. Both models are derived through supervised fine-tuning on a carefully curated dataset of 800k examples distilled from DeepSeek-R1 [guo2025deepseek]. The respective base models used for distillation are Qwen2.5-Math-7B and Llama-3.1-8B [yang2024qwen2, grattafiori2024llama]. Through the distillation process, these models inherit the CoT reasoning capabilities of DeepSeek-R1. During inference, their reasoning steps are explicitly structured and enclosed within <think></think> tags. For clarity and brevity throughout the paper, we refer to these models as R1-Qwen-7B and R1-Llama-8B, respectively.
Comparison methods. In addition to the vanilla model, we select two comparative methods: In-context Alignment (ICA) [lin2023unlocking] and Self-Align [sun2023principle], each implemented with two variants. ICA is a context-based alignment approach that does not require fine-tuning. Specifically, we use GPT-4o to generate responses for samples from the training set. The first variant, ICA (Random), randomly selects 1 harmful and 1 benign (prompt, response) pairs from the training set as in-context examples. For the second variant, ICA (Curated), we manually select 1 harmful and 1 benign correct samples from the GPT-4o-generated responses to construct the in-context examples. Self-Align is a fine-tuning-based alignment method in which a model samples its own responses and ranks them to identify high-quality outputs, which are then used as synthetic data to further train the model. In this setup, a vanilla model is used for both response generation and selection. Specifically, for Self-Align (SFT), the response with the highest score is selected as the target to construct the fine-tuning dataset. For Self-Align (DPO), the highest- and lowest-scoring responses are treated as the chosen and rejected samples, respectively. Both variants rely on fine-tuning the vanilla model.
@@
Evaluation. Following the work of Guan et al. [guan2024deliberative], we evaluate the model’s performance in terms of both safety against harmful prompts and its ability to avoid excessive refusals of benign prompts. These aspects are quantified using the “Not unsafe” and “Not overrefuse” metrics, respectively. The “Not unsafe” metric assesses the model’s capacity to maintain safety when confronted with harmful inputs. It is defined as the proportion of harmful samples that GPT-4o labels as safe, relative to the total number of harmful samples. The “Not overrefuse” metric measures the model’s tendency to avoid unnecessary refusals when dealing with benign inputs. It is computed as the proportion of benign samples that GPT-4o judges as not excessively refused, out of the total number of benign samples. An example of the evaluation instruction is provided in Figure 6. To obtain an overall performance score, we average the values of the two metrics.
@@
For training, we adopt the LoRA [Hu2021LoRALA] efficient parameter fine-tuning method. All experiments are conducted over a single epoch, with a learning rate of 5e-4, a warm-up phase comprising 10% of the total training steps, and a cosine learning rate scheduler. For inference, we use the vLLM [kwon2023efficient] efficient inference framework, setting the temperature to 0.6 and the maximum sequence length to 2048, while keeping all other parameters at their default values. All experiments are run on a single 80GB A800 GPU.
@@ 4.3 Impact of Data Quantity and Distribution on Alignment Fine-Tuning
In this subsection, we examine how the quantity and ratio of synthetic data, generated according to safety guidelines, influence model performance during the alignment process. Firstly, as shown in Figure 7 (left), we present experimental results analyzing changes in performance during the SFT stage across three datasets. Notably, the ratio of harmful to benign data is controlled at 1:1. We highlight three key findings: (1) Performance improvement with diminishing returns: Increasing the amount of training data generally enhances overall model performance, though marginal gains decrease as the volume grows. Specifically, when scaling from 200 to 800 samples, models exhibit improved safety scores, albeit with a slight increase in overrefuse. Nonetheless, the overall alignment performance improves. However, at 1000 samples, performance begins to slightly decline across multiple models and datasets, suggesting the existence of the risk of overfitting. This may be attributed to the model’s inherent performance ceiling as well as unavoidable noise introduced by self-synthesized data. Therefore, when leveraging self-synthesized data for alignment enhancement, it is important to balance the benefits of increased scale against potential risks. On one hand, the data augmentation advantage should be fully utilized; on the other hand, stricter quality filtering mechanisms should be considered to mitigate the risk of bias introduced by feedback loops in the model [liu2024best, dong2025selfboosting]. (2) The stronger model exhibits better self-alignment abilities: R1-Llama-8B demonstrates superior stability and alignment efficiency compared to R1-Qwen-7B. Across all experiments, R1-Llama-8B consistently outperforms R1-Qwen-7B across the three evaluation metrics. This advantage can be attributed to the larger parameter scale of R1-Llama-8B, which enables more effective synthetic data generation and indicates a stronger capability to learn meaningful patterns from synthetic alignment data. (3) Task-specific sensitivity to data quantity: The effect of synthetic data volume on SFT outcomes varies by task. We observe that the performance of the MaliciousEducator and WildJailbreak datasets remains relatively stable with increased data, while MathPrompt shows more volatility. This indicates that the optimal amount of fine-tuning data should be determined with consideration of the specific characteristics of the target task. For instance, adversarial jailbreak like MathPrompt may not benefit linearly from simply increasing the volume of data during the SFT phase. Instead, alternative alignment methods such as DPO, which focus on distinguishing between positive and negative samples, may be more effective for improving model safety alignment in a more targeted manner.
@@
The results of the data ratio experiments conducted during the DPO stage are presented in Figure 7 (right), from which we make the following observations: First, we find an intuitive trade-off: increasing the proportion of benign prompts generally leads to a decline in model safety, while simultaneously reducing the degree of overrefuse. For instance, in the experiment using the R1-Llama-8B model on the WildJailbreak dataset, the safety score decreased from 0.914 to 0.876, whereas the “Not overrefuse” score improved from 0.843 to 0.886. This trade-off has been widely reported in recent literature and is considered an inherent challenge in the alignment process [tan2025equilibraterlhfbalancinghelpfulnesssafety]. Second, we observe that tuning the data ratio plays a crucial role in achieving fine-grained behavioral alignment. A 1:1 ratio is often suboptimal; alternative configurations such as 3:5 or 5:3 typically yield higher average scores. This suggests that moderately skewed distributions—favoring either benign or harmful samples—can better support the desired alignment behavior, and should be selected based on the model’s characteristics following supervised fine-tuning (SFT). In summary, the distribution of training samples has a significant impact on the final alignment equilibrium of the model. Adjusting the ratio of harmful to benign prompts serves as an effective means of setting behavioral thresholds and delineating the boundaries of the model’s responses.

### v2

@@ 1 Introduction
Reasoning models, including ChatGPT-o1 [1], ChatGPT-o3 [2], and DeepSeek-R1 [3], have demonstrated remarkable capabilities across diverse and complex reasoning tasks, such as code generation, mathematical problem-solving, multi-hop question answering, and decision-making [4, 5, 6]. A pivotal factor behind this success is the Chain-of-Thought (CoT) prompting strategy [7], which guides models to explicitly generate intermediate reasoning steps before concluding with the final answer. This approach significantly enhances the accuracy and interpretability of the model’s outputs by effectively leveraging its intrinsic reasoning capabilities. Beyond achieving high performance, ensuring the safety of reasoning models is increasingly critical [8, 9]. As these models become more prevalent in real-world applications, it is essential to prevent them from generating harmful, unethical, or adversarial content.
However, the safety mechanisms acquired by reasoning models through alignment algorithms [10, 11] demonstrate certain effectiveness in handling standard (vanilla) jailbreak prompts, but they still exhibit vulnerabilities when confronted with more complex adversarial jailbreak techniques. Figure 1, based on the MaliciousEducator dataset [12], illustrates examples of vanilla harmful and benign prompts, along with their adversarial variants. Notably, adversarial harmful prompts frequently conceal malicious intent through indirect, covert strategies, such as framing harmful requests as educational inquiries. Conversely, adversarial benign prompts structurally resemble jailbreak attempts but do not harbor malicious intent. As demonstrated in the model responses shown in Figure 1, the model should be capable of refusing responses to adversarial harmful prompts, while avoiding excessive rejection of adversarial benign prompts.
Additionally, adversarial jailbreak prompts generally exhibit two distinctive features: (1) Limited availability, which includes a small number of prompts (usually only dozens) and an almost complete absence of corresponding reasoning responses suitable for fine-tuning. (2) High structural and semantic variability, such as mathematical encodings and pedagogically styled questions [12, 13, 14]. These characteristics pose significant difficulties for traditional alignment methods reliant on extensive human intervention, limiting their scalability and responsiveness to rapidly evolving threats.
@@
We detail our methodology further in Section 3. Extensive experiments across two reasoning models and three datasets demonstrate our framework’s efficacy. Notably, on the MaliciousEducator dataset, SGASA achieved improvements of 34.5% and 32.2% using DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B [3] as backbones, respectively. We also explore the impact of data volume and ratio on alignment fine-tuning results and perform cross-dataset evaluations, verifying our approach’s generalization across different attack strategies. Detailed discussions and findings are presented in Section 4.
@@ 2 Related Work
Ensuring the safety alignment of large language models has been a key research direction to guarantee their reliability and prevent the generation of harmful content. A substantial body of work has explored various alignment methods. For instance, RLHF [10] trains a reward model based on human rankings of multiple model outputs, and then fine-tunes the language model using this reward signal to produce responses preferred by humans. DPO [11] bypasses the separate reward-model training step by converting human comparison data directly into a loss on the model’s own policy. Beyond these, with the improvement of model reasoning capabilities – especially the emergence of large reasoning models such as OpenAI’s o1 series [1], DeepSeek-R1 [15], Alibaba’s Qwen model [16], Google’s Gemini Flash Thinking model [17], and the KIMI reasoning model [18] – a set of “slow thinking” alignment techniques has been proposed to enhance model safety. In this line of work, researchers directly teach the model textual safety specifications and train it to explicitly recall and reason over these rules before generating an answer. For example, deliberative alignment [19] teaches the model to apply written policy rules via Chain-of-Thought reasoning. Applied to OpenAI’s o‑series models, deliberative alignment demonstrated that models can employ Chain‑of‑Thought reasoning to identify relevant policy clauses and produce safer, more compliant responses, achieving Pareto improvements by reducing over‑refusal while boosting adversarial robustness. Similarly, SafeChain [20] fine-tunes models on CoT–style safety data. Other approaches address related issues, such as reducing undue refusals through multi-turn reasoning [21] or using introspective CoT analysis to identify safety risks [22].
@@
Many studies focus on how to bypass a model’s safety alignment mechanisms to induce it to generate harmful content, using approaches such as decoding strategies, prefilling strategies [23, 24], and optimization strategies [25, 26, 27]. Another line of work hides harmful intent within specific contextual forms. This approach misleads the model with seemingly benign but actually malicious inputs, causing it to output unsafe information. For example, [14] transforms harmful user requests into mathematical symbols, while [12] disguises them as educational scenarios (such as law enforcement training). The approach proposed in this paper mainly targets this type of jailbreak attack conducted through natural language.
@@
Given the rapid advances in the capabilities of Large Language Models (LLMs), traditional alignment approaches reliant on extensive human annotation have become increasingly difficult to scale. Consequently, a range of automated alignment methods have been extensively studied to mitigate these scalability limitations. A notable category among these methods is model self-alignment. Sun et al. (2023) [28] proposed principle-driven self-alignment, aiming to explicitly guide models according to a concise set of human-defined principles, thus significantly minimizing the need for human supervision. Zhang et al. (2024) [29] introduced a self-alignment strategy designed to reduce factual hallucinations by exploiting the intrinsic self-evaluation capabilities of language models. Their proposed approach, consisting of a self-evaluation mechanism combined with Self-Knowledge Tuning, demonstrated considerable improvements in factual accuracy without the need for external human annotations. Furthermore, Xu et al. (2025)  [30] presented PT-ALIGN, a dual safety self-alignment method targeting both positive and toxic samples. By integrating Maximum Likelihood Estimation for positive examples and fine-grained Unlikelihood Training for toxic ones, PT-ALIGN effectively enhances the safety of model outputs while substantially reducing human intervention. Overall, the research on self-alignment methods has shown significant potential in reducing human involvement and improving output quality, establishing itself as an increasingly critical direction within the broader field of language model alignment [31].
@@ 4.1 Experimental Settings
  • WildJailbreak [13] is an open-source safety dataset whose test set includes two types of prompts: adversarial harmful prompts, which convey harmful requests in more convoluted and stealthy ways, and adversarial benign prompts, which are similar in form but do not carry harmful intent. In our study, we construct a training subset by sampling five prompts of each type from the original training set. For evaluation, we randomly select 210 prompts of each type to create a test set comprising a total of 420 prompts.
  • MathPrompt [14] proposes a novel jailbreaking technique that reformulates harmful user requests into symbolic mathematical expressions to bypass safety mechanisms. Since the original dataset used in the study is not publicly available, we reconstruct the test set based on the described methodology. For the training set, we include five prompts of each type. The testing set consists of 150 harmful prompts and 150 benign prompts.
  • MaliciousEducator[12] is a jailbreak dataset developed for educational purposes. It contains 50 samples featuring highly dangerous queries and suspicious intents. Based on this dataset, we add an equal number of benign prompts with similar structure and format. We randomly select 5 samples of each type to form the training set. The resulting test set consists of 45 harmful samples and 45 benign samples.
@@
Large Reasoning Models. In this work, we investigate two reasoning models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B. Both models are derived through supervised fine-tuning on a carefully curated dataset of 800k examples distilled from DeepSeek-R1 [3]. The respective base models used for distillation are Qwen2.5-Math-7B and Llama-3.1-8B [32, 33]. Through the distillation process, these models inherit the CoT reasoning capabilities of DeepSeek-R1. During inference, their reasoning steps are explicitly structured and enclosed within <think></think> tags. For clarity and brevity throughout the paper, we refer to these models as R1-Qwen-7B and R1-Llama-8B, respectively.
Comparison methods. In addition to the vanilla model, we select two comparative methods: In-context Alignment (ICA) [34] and Self-Align [28], each implemented with two variants. ICA is a context-based alignment approach that does not require fine-tuning. Specifically, we use GPT-4o to generate responses for samples from the training set. The first variant, ICA (Random), randomly selects 1 harmful and 1 benign (prompt, response) pairs from the training set as in-context examples. For the second variant, ICA (Curated), we manually select 1 harmful and 1 benign correct samples from the GPT-4o-generated responses to construct the in-context examples. Self-Align is a fine-tuning-based alignment method in which a model samples its own responses and ranks them to identify high-quality outputs, which are then used as synthetic data to further train the model. In this setup, a vanilla model is used for both response generation and selection. Specifically, for Self-Align (SFT), the response with the highest score is selected as the target to construct the fine-tuning dataset. For Self-Align (DPO), the highest- and lowest-scoring responses are treated as the chosen and rejected samples, respectively. Both variants rely on fine-tuning the vanilla model.
@@
Evaluation. Following the work of Guan et al. [19], we evaluate the model’s performance in terms of both safety against harmful prompts and its ability to avoid excessive refusals of benign prompts. These aspects are quantified using the “Not unsafe” and “Not overrefuse” metrics, respectively. The “Not unsafe” metric assesses the model’s capacity to maintain safety when confronted with harmful inputs. It is defined as the proportion of harmful samples that GPT-4o labels as safe, relative to the total number of harmful samples. The “Not overrefuse” metric measures the model’s tendency to avoid unnecessary refusals when dealing with benign inputs. It is computed as the proportion of benign samples that GPT-4o judges as not excessively refused, out of the total number of benign samples. An example of the evaluation instruction is provided in Figure 6. To obtain an overall performance score, we average the values of the two metrics.
@@
For training, we adopt the LoRA [35] efficient parameter fine-tuning method. All experiments are conducted over a single epoch, with a learning rate of 5e-4, a warm-up phase comprising 10% of the total training steps, and a cosine learning rate scheduler. For inference, we use the vLLM [36] efficient inference framework, setting the temperature to 0.6 and the maximum sequence length to 2048, while keeping all other parameters at their default values. All experiments are run on a single 80GB A800 GPU.
@@ 4.3 Impact of Data Quantity and Distribution on Alignment Fine-Tuning
In this subsection, we examine how the quantity and ratio of synthetic data, generated according to safety guidelines, influence model performance during the alignment process. Firstly, as shown in Figure 7 (left), we present experimental results analyzing changes in performance during the SFT stage across three datasets. Notably, the ratio of harmful to benign data is controlled at 1:1. We highlight three key findings: (1) Performance improvement with diminishing returns: Increasing the amount of training data generally enhances overall model performance, though marginal gains decrease as the volume grows. Specifically, when scaling from 200 to 800 samples, models exhibit improved safety scores, albeit with a slight increase in overrefuse. Nonetheless, the overall alignment performance improves. However, at 1000 samples, performance begins to slightly decline across multiple models and datasets, suggesting the existence of the risk of overfitting. This may be attributed to the model’s inherent performance ceiling as well as unavoidable noise introduced by self-synthesized data. Therefore, when leveraging self-synthesized data for alignment enhancement, it is important to balance the benefits of increased scale against potential risks. On one hand, the data augmentation advantage should be fully utilized; on the other hand, stricter quality filtering mechanisms should be considered to mitigate the risk of bias introduced by feedback loops in the model [37, 38]. (2) The stronger model exhibits better self-alignment abilities: R1-Llama-8B demonstrates superior stability and alignment efficiency compared to R1-Qwen-7B. Across all experiments, R1-Llama-8B consistently outperforms R1-Qwen-7B across the three evaluation metrics. This advantage can be attributed to the larger parameter scale of R1-Llama-8B, which enables more effective synthetic data generation and indicates a stronger capability to learn meaningful patterns from synthetic alignment data. (3) Task-specific sensitivity to data quantity: The effect of synthetic data volume on SFT outcomes varies by task. We observe that the performance of the MaliciousEducator and WildJailbreak datasets remains relatively stable with increased data, while MathPrompt shows more volatility. This indicates that the optimal amount of fine-tuning data should be determined with consideration of the specific characteristics of the target task. For instance, adversarial jailbreak like MathPrompt may not benefit linearly from simply increasing the volume of data during the SFT phase. Instead, alternative alignment methods such as DPO, which focus on distinguishing between positive and negative samples, may be more effective for improving model safety alignment in a more targeted manner.
@@
The results of the data ratio experiments conducted during the DPO stage are presented in Figure 7 (right), from which we make the following observations: First, we find an intuitive trade-off: increasing the proportion of benign prompts generally leads to a decline in model safety, while simultaneously reducing the degree of overrefuse. For instance, in the experiment using the R1-Llama-8B model on the WildJailbreak dataset, the safety score decreased from 0.914 to 0.876, whereas the “Not overrefuse” score improved from 0.843 to 0.886. This trade-off has been widely reported in recent literature and is considered an inherent challenge in the alignment process [39]. Second, we observe that tuning the data ratio plays a crucial role in achieving fine-grained behavioral alignment. A 1:1 ratio is often suboptimal; alternative configurations such as 3:5 or 5:3 typically yield higher average scores. This suggests that moderately skewed distributions—favoring either benign or harmful samples—can better support the desired alignment behavior, and should be selected based on the model’s characteristics following supervised fine-tuning (SFT). In summary, the distribution of training samples has a significant impact on the final alignment equilibrium of the model. Adjusting the ratio of harmful to benign prompts serves as an effective means of setting behavioral thresholds and delineating the boundaries of the model’s responses.